# INTRABENCH: Interactive Radiological Benchmark

### v1

@@ A.1 SAM
SAM is compatible with multiple image encoders, particularly the ViT family from Dosovitskiy et al. (2021). We used the default and best-performing model with ViT-Huge. To ensure high-quality inputs for the model, we performed slice-wise inference by extracting slices from the inplane-plane axis. Each slice was normalized by first clipping values outside the 0.5th and 99.5th percentile of the volume’s intensity distribution and then scaling the values to [0, 255]. The image was repeated three times along the channel axis to produce an RGB-like image. Internally, SAM resizes these slices to 1024 pixels for the longest side with the shorter side being padded to 1024 pixels if needed to maintain square dimensions. Finally, the images are normalized using the model’s pre-stored mean and standard deviation as suggested by the original implementation. Inference was restricted to slices containing foreground. After prediction, the slices were reassembled into a volume, inverse transformed to the original coordinate system, and metrics were computed in the original image space.
A.2 SAM2
@@
SAM2 supports multiple image encoders, specifically the Hiera family of Ryali et al. (2023). We used the best-performing model, Hiera-L. We clip the intensity values of the volumes based on the 0.5th and 99.5th percentiles, extract each slice along the through-plane, and make the images RGB-like just as with SAM. The images are then rescaled to $1024\times 1024$ pixels and again normalized using the mean and standard deviation provided together with the pretrained weights. Aggregation and inverse transformation are then performed similarly to SAM.
A.3 MedSAM
@@ A.4 SAM-Med2D
A.5 SAM-Med3D
The model is 3D so no slicing is needed. The volume is respaced to $1.5\times 1.5\times 1.5$ mm and then clipped based on its 0.5th and 99.5th percentiles. SAMMed3D performs inference on a 128x128x128 crop. The crop is centered around our point prompt if there is only one point prompt passed, and around the centroid of our prompts if multiple points are passed simultaneously. For subsequent refinement steps, the crop remains unchanged. The predicted crop is inserted back in its correct position within the wider coordinate system and then respaced back to the original spacing so that evaluation takes place in the corresponding native image space.
A.6 SAM-Med3D Turbo
SAM-Med3D Turbo is an updated checkpoint for SAMMed-3D and so we perform the same pre- and postprocessing.
A.7 SegVol
@@
Appendix B Additional Results
B.1 Initial Prediction - Unrealistic Effort
@@
Table 2: Experimental results simulating unrealistic effort of a clinician prompting each slice of a 3D volume. ’PPS’ and ’BPS’ represent points per slice or box per slice, respectively. ’X’ implies that each interaction is replicated for every slice, multiplying the clinician’s effort across the entire volume.
B.2 Single Forward Pass - Realistic Effort 2D
@@ B.3 Single Forward Pass - Realistic Effort 3D
Table 4: Experimental results simulating a realistic clinician’s effort. ’center PPV’ stands for Point Per Volume, that was sampled from the center of the target object. Sampling from the center performs better than sampling a random point. *SegVol used D2 HanSeg as training data.
B.4 Iterative Refinement 2D
@@
0 3B Inter + Scribble Refine SAM 6/3 13.13 54.04 63.34 58.64 64.26 71.35 65.52 61.78 25.58 59.23 53.69
1 3B Inter + Scribble Refine SAM 6/3 0.99 3.66 3.24 3.53 2.33 6.14 7.78 11.88 4.68 6.43 5.07
2 3B Inter + Scribble Refine SAM 6/3 0.85 1.95 1.28 3.17 9.04 5.28 7.49 14.4 4.14 7.43 5.51
3 3B Inter + Scribble Refine SAM 6/3 0.8 4.11 2.95 2.19 3.13 3.36 7.14 9.85 4.02 4.67 4.22
4 3B Inter + Scribble Refine SAM 6/3 0.71 2.45 2.01 2.21 2.28 1.89 6.08 9.09 3.22 3.57 3.35
5 3B Inter + Scribble Refine SAM 6/3 0.76 1.99 2.46 2.51 2.16 1.42 6.15 9.34 3.19 3.47 3.34
0 3B Inter + Scribble Refine SAM2 6/3 69.78 57.59 67.71 69.08 69.06 73.11 68.35 63.24 37.21 63.27 63.84
1 3B Inter + Scribble Refine SAM2 6/3 10.03 20.17 25.37 19.13 23.65 18.12 25.37 27.66 11.34 15.93 19.68
2 3B Inter + Scribble Refine SAM2 6/3 5.07 6.98 10.53 7.31 8.35 3.99 20.7 19.58 5.61 6.41 9.45
3 3B Inter + Scribble Refine SAM2 6/3 4.23 4.43 7.8 6.93 5.64 3.72 14.43 16.35 4.89 5.34 7.38
4 3B Inter + Scribble Refine SAM2 6/3 3.99 3.84 6.96 7.22 4.86 3.53 13.37 15.02 4.56 4.9 6.83
5 3B Inter + Scribble Refine SAM2 6/3 3.41 3.7 6.75 5.7 3.99 2.79 12.49 13.59 4.3 4.47 6.12
0 3B Inter + Scribble Refine SamMed 2D 6/3 27.83 42.22 52.7 60.66 44.76 38.63 55.31 55.81 33.46 38.66 45.0
1 3B Inter + Scribble Refine SamMed 2D 6/3 28.4 44.72 57.53 63.33 45.28 41.42 61.61 59.73 40.44 38.8 48.13
2 3B Inter + Scribble Refine SamMed 2D 6/3 24.29 42.41 53.02 59.24 40.27 35.69 62.86 58.85 41.23 34.93 45.28
3 3B Inter + Scribble Refine SamMed 2D 6/3 20.91 39.69 47.48 52.73 35.55 28.2 62.08 54.89 40.02 31.59 41.31
4 3B Inter + Scribble Refine SamMed 2D 6/3 18.76 37.29 41.46 46.5 31.73 23.98 58.46 51.86 38.07 28.74 37.69
5 3B Inter + Scribble Refine SamMed 2D 6/3 17.07 35.06 37.39 41.62 29.1 20.8 53.9 49.27 36.27 26.66 34.71
@@
Table 5: Interactive refinement results for 2D models across 5 iterations. The initial prediction is made either using a single point per slice or one of our proposed prompting schemes. Omitting the previous point during refinement led to a drop in performance, which explains the drop in performance for the 3 Box interpolation after the initial prompt.The unrealistic slice-wise refinement (1 interaction per slice) is only slightly better than our proposed scribble refinement method (3 interactions).
B.5 Iterative Refinement 3D
@@
0 1 center PPV + 1 PPV Refine + prev point SamMed 3D 1/1 2.03 12.15 24.11 26.9 15.09 19.64 72.66 53.03 50.4 26.51 30.25
1 1 center PPV + 1 PPV Refine + prev point SamMed 3D 1/1 5.89 14.07 33.03 35.67 14.42 24.93 64.15 72.35 42.64 24.86 33.2
2 1 center PPV + 1 PPV Refine + prev point SamMed 3D 1/1 5.67 13.49 29.52 28.92 14.34 21.4 64.29 62.44 42.03 25.16 30.73
3 1 center PPV + 1 PPV Refine + prev point SamMed 3D 1/1 5.57 13.64 31.57 35.51 14.21 22.82 63.35 65.94 41.12 24.79 31.85
4 1 center PPV + 1 PPV Refine + prev point SamMed 3D 1/1 5.81 13.51 31.44 31.08 14.11 21.38 65.28 63.38 40.68 24.82 31.15
5 1 center PPV + 1 PPV Refine + prev point SamMed 3D 1/1 5.46 13.56 32.19 33.49 14.18 24.28 62.12 65.2 40.41 25.12 31.6
0 1 center PPV + 1 PPV Refine + prev point SamMed 3D Turbo 1/1 5.12 27.34 46.07 34.46 15.9 46.38 82.98 59.36 63.83 26.98 40.84
1 1 center PPV + 1 PPV Refine + prev point SamMed 3D Turbo 1/1 2.81 21.83 27.6 17.69 8.25 30.88 79.91 72.82 61.07 21.37 34.42
2 1 center PPV + 1 PPV Refine + prev point SamMed 3D Turbo 1/1 3.57 22.72 35.57 22.05 8.89 36.31 77.77 72.2 62.7 23.37 36.52
3 1 center PPV + 1 PPV Refine + prev point SamMed 3D Turbo 1/1 3.25 22.22 32.87 17.88 8.74 33.27 77.73 66.93 61.51 22.17 34.66
4 1 center PPV + 1 PPV Refine + prev point SamMed 3D Turbo 1/1 3.6 22.93 34.4 16.26 8.78 34.84 80.01 72.09 62.35 23.35 35.86
5 1 center PPV + 1 PPV Refine + prev point SamMed 3D Turbo 1/1 3.09 21.72 32.32 14.83 7.66 31.79 76.55 70.69 62.1 22.19 34.3
@@
Table 6: Interactive refinement results for 3D models over 5 iterations. The initial interaction always starts from a central point of the target object, and refinement is performed either by randomly sampling positive or negative points (1 interaction) or by selecting a point using the proposed scribble refinement method. Scribble drawing is counted as three interactions. In contrast to 2D models, including the previous point prompt did not improve the performance.

### v2

@@ 5.1 SAM
SAM is compatible with multiple image encoders, particularly the ViT family from [8]. We used the default and best-performing model with ViT-Huge. To ensure high-quality inputs for the model, we performed slice-wise inference by extracting slices from the inplane-plane axis. Each slice was normalized by first clipping values outside the 0.5th and 99.5th percentile of the volume’s intensity distribution and then scaling the values to [0, 255]. The image was repeated three times along the channel axis to produce an RGB-like image. Internally, SAM resizes these slices to 1024 pixels for the longest side with the shorter side being padded to 1024 pixels if needed to maintain square dimensions. Finally, the images are normalized using the model’s pre-stored mean and standard deviation as suggested by the original implementation. Inference was restricted to slices containing foreground. After prediction, the slices were reassembled into a volume, inverse transformed to the original coordinate system, and metrics were computed in the original image space.
5.2 SAM2
@@
SAM2 supports multiple image encoders, specifically the Hiera family of Ryali et al. [35]. We used the best-performing model, Hiera-L. We clip the intensity values of the volumes based on the 0.5th and 99.5th percentiles, extract each slice along the through-plane, and make the images RGB-like just as with SAM. The images are then rescaled to $1024\times 1024$ pixels and again normalized using the mean and standard deviation provided together with the pretrained weights. Aggregation and inverse transformation are then performed similarly to SAM.
5.3 MedSAM
@@ 5.4 SAM-Med2D
5.5 SAM-Med3D
The model is 3D so no slicing is needed. The volume is respaced to $1.5\times 1.5\times 1.5$ mm and then clipped based on its 0.5th and 99.5th percentiles. SAM-Med 3D performs inference on a 128x128x128 crop. The crop is centered around our point prompt if there is only one point prompt passed, and around the centroid of our prompts if multiple points are passed simultaneously. For subsequent refinement steps, the crop remains unchanged. The predicted crop is inserted back in its correct position within the wider coordinate system and then respaced back to the original spacing so that evaluation takes place in the corresponding native image space.
5.6 SAM-Med3D Turbo
SAM-Med3D Turbo is an updated checkpoint for SAM-Med-3D and so we perform the same pre- and postprocessing.
5.7 SegVol
@@
6 Supported prompting schemes
2D initial prompts:
For the 2D models, every foreground-containing slice needs an initial prompt. RadioActive supports the following 2D prompting schemes:
  • N PPS: For each foreground-containing slice, N random foreground points are selected.
  • N±$\pm$±PPS: For each foreground-containing slice, $\lceil N/2\rceil$ random foreground and $\lfloor N/2\rfloor$ background points are selected.
  • Box PS: For each foreground-containing slice, we select a bounding box around the foreground.
  • NP Inter: N points are used for the proposed point interpolation prompting scheme as explained in Sec. 2.1. At least 3 points are needed.
  • P Prop: The center point of the slice with index $I_{m}$ is selected, and passed to the model along with $min(I)$ and $max(I)$ with notation and method as explained in Sec. 2.1.
  • NB Inter: N 2D boxes are used for the proposed box interpolation prompting scheme as explained in Sec. 2.1. At least 3 boxes are needed.
  • B Prop: The 2D box from the slice with index $I_{m}$ is selected, and passed to the model along with $min(I)$ and $max(I)$ with notation and method as explained inSec. 2.1.
3D initial prompts
For 3D models, the entire volume is predicted in a single step, making the prompting process more accessible for human interaction.
  • N PPV: N random foreground points are selected from the volume.
  • N center PPV: We generate a set of slice points using the point interpolation method as in Sec. 2.1 and then select a subset o $N$ points spaced equally along the axial direction.
  • 3D Box: A simple 3D bounding box around the foreground is used.
2D iterative refinement
The refinement experiments always consist of an initial prompt, followed by iterative refinement prompts that are generated as explained in the following. For all refinement steps, a point prompt from a previous step can be included.
  • 1PPS Refine: For each slice, a refinement point prompt is selected randomly from the misclassified pixels, and passed either a false positive or false negative as appropriate from the false positive or false negative predicted volume.
  • Scribble Refine: This prompting scheme follows the proposed scribble refinement scheme for 2D models in Sec. 2.2. Only slices that receive a new refinement prompt are predicted again.
3D iterative refinement
Again, a point prompt from a previous step can be reused in all subsequent refinement steps.
  • 1 PPV Refine: For each refinement step, one positive or negative point prompt per volume is selected randomly from the misclassified pixels, and passed either a false positive or false negative as appropriate.
  • Scribble Refine: We generate a refinement scribble as we would for 2D iterative refinement, but then randomly select just one refinement point from the scribble.
All 2D and 3D bounding boxes were generated with perfect alignment around the target object. Since the impact of non-perfect bounding boxes has been extensively studied in previous works, we chose not to explore this additional degree of freedom [34, 27]. However, the open-source nature of our framework facilitates straightforward extensions to include a wider range of prompting scenarios.
7 Additional Results
* Refer to caption
Figure 8: Qualitative examples of failure-cases: A) We highlight issues of SAM-Med3D, which is limited to predict within a small field-of-view. Segmentations are truncated at the edges, due to the organ exceeding the boundaries of the prediction volume. B) 2D segmentation models sometimes default to predicting the entire abdomen, due to the ambiguity of what was queried.
Difference between MRI und CT data
Table 8 presents the DSC for prompting schemes simulating realistic Human Effort without refinement, averaged over MRI datasets (D1-3) and CT datasets (D4-10). While direct comparisons between modalities are not possible due to differing tasks and datasets, notable trends emerge. Some models exhibit consistent performance across modalities, such as SAM2, which shows minimal variation. In contrast, all 3D models display significant performance differences between MRI and CT. This disparity likely reflects the training data distribution; for instance, SegVol was exclusively trained on CT data, leading to subpar performance on MRI datasets. However, on CT-specific tasks, SegVol matches the performance of SAM and SAM2, despite its overall lower average performance across all datasets.
Previous prompts are essential for 2D model refinement
For 2D models, it is important to reintroduce the initial prompts at each refinement step. In Tab. 5, we demonstrate that omitting the initial prompt during iterative refinement results in a performance decline. This behavior arises because 2D models often over-segment the target, filling the entire slice foreground. This is visualized in Fig. 8 B. Without the initial prompt, these models lose crucial target location information, as the initial predicted mask is typically highly inaccurate. Consequently, the refinement process likely generates erroneous additional prompts due to the abundance of false-positive pixels, further degrading performance.
Previous prompts hinder 3D model refinement
For 3D models, Tab. 6 demonstrates that incorporating previous prompts does not enhance refinement performance beyond the first step. This observation aligns with the findings in Fig. 6 (right), where SAM-Med 3D and SAM-Med 3D Turbo show no improvement when provided with more than three prompts simultaneously.
Performance comparison between organs and pathological structures
Table 7 presents a performance comparison between pathological and organ targets. While a direct comparison is challenging due to variations in tasks and datasets, certain trends can be observed. Notably, 2D models exhibit greater difficulty in segmenting pathologies when prompted with points alone, likely due to the typically higher contrast of organs compared to pathological structures.
Challenges of 3D patch-based prediction for larger structures
Table 7 shows the DSC for liver segmentation (D8a), a common task included in the training datasets of all medical models. However, SAM-Med 3D and SAM-Med 3D Turbo encounter difficulties in predicting the entire liver within a single patch. In contrast, SegVol’s zoom-out/zoom-in inference approach effectively addresses this limitation, achieving a superior DSC of 91%. Such a failure-case is visualized in Fig. 8 A.
@@ 7 Additional Results
Table 2: Experimental results simulating unrealistic effort of a clinician. This involves prompting each slice of a 3D volume. ’PPS’ and ’BPS’ represent points per slice or box per slice, respectively. ’X’ implies that each interaction is replicated for every slice, multiplying the clinician’s effort across the entire volume.
@@
Table 4: Experimental results simulating a realistic clinician’s effort. ’center PPV’ stands for Point Per Volume, that was sampled from the center of the target object. Sampling from the center performs better than sampling a random point. SegVol trained on CT images from the D2 HanSeg dataset, while we tested on MRI images from the same patient cohort.
@@
0 1PPS + 1PPS Refine * SAM 1X/1X 0.81 2.66 3.13 2.06 1.42 1.0 7.86 12.47 3.42 3.33 3.82
1 1PPS + 1PPS Refine * SAM 1X/1X 0.06 0.27 0.63 0.6 0.15 0.08 3.34 3.38 1.45 0.61 1.06
2 1PPS + 1PPS Refine * SAM 1X/1X 0.33 0.95 1.55 1.64 0.76 0.67 6.2 11.11 2.61 2.45 2.83
3 1PPS + 1PPS Refine * SAM 1X/1X 0.22 1.06 1.56 1.42 0.86 0.51 5.16 7.74 2.14 1.54 2.22
4 1PPS + 1PPS Refine * SAM 1X/1X 0.38 1.08 1.73 1.4 0.5 0.57 6.07 10.85 2.61 2.3 2.75
5 1PPS + 1PPS Refine * SAM 1X/1X 0.2 0.93 1.34 1.29 0.62 0.58 5.53 8.34 2.07 1.69 2.26
0 1PPS + 1PPS Refine * SAM2 1X/1X 1.25 2.74 4.08 3.9 3.07 1.51 9.6 16.35 3.97 3.83 5.03
1 1PPS + 1PPS Refine * SAM2 1X/1X 1.33 2.94 5.86 4.34 3.44 1.76 11.73 18.45 4.13 3.74 5.77
2 1PPS + 1PPS Refine * SAM2 1X/1X 1.17 3.02 6.24 4.62 3.53 1.64 11.58 17.89 4.3 3.66 5.77
3 1PPS + 1PPS Refine * SAM2 1X/1X 1.29 3.05 6.31 4.93 3.69 1.73 11.78 17.81 4.34 3.67 5.86
4 1PPS + 1PPS Refine * SAM2 1X/1X 1.31 3.22 6.57 4.97 3.79 1.75 11.82 17.5 4.37 3.84 5.91
5 1PPS + 1PPS Refine * SAM2 1X/1X 1.23 3.21 6.6 4.86 3.74 1.76 11.97 17.78 4.4 3.8 5.93
0 1PPS + 1PPS Refine * SamMed 2D 1X/1X 10.72 26.3 24.55 28.83 22.76 9.71 30.83 31.24 26.05 20.31 23.13
1 1PPS + 1PPS Refine * SamMed 2D 1X/1X 10.45 28.62 27.4 31.37 23.26 10.47 38.52 37.22 30.63 20.04 25.8
2 1PPS + 1PPS Refine * SamMed 2D 1X/1X 9.88 28.71 28.05 31.7 22.67 10.37 41.79 38.9 31.55 19.47 26.31
3 1PPS + 1PPS Refine * SamMed 2D 1X/1X 9.6 28.45 28.26 31.42 22.18 10.58 43.5 40.16 31.9 18.97 26.5
4 1PPS + 1PPS Refine * SamMed 2D 1X/1X 9.2 28.17 28.29 31.08 21.78 10.44 44.71 40.81 31.99 18.52 26.5
5 1PPS + 1PPS Refine * SamMed 2D 1X/1X 8.99 27.84 28.26 30.56 21.49 10.54 45.45 41.43 31.99 18.2 26.48
@@
0 3B Inter + Scribble Refine* SAM 6/3 13.13 54.04 63.34 58.64 64.26 71.35 65.52 61.78 25.58 59.23 53.69
1 3B Inter + Scribble Refine* SAM 6/3 0.99 3.66 3.24 3.53 2.33 6.14 7.78 11.88 4.68 6.43 5.07
2 3B Inter + Scribble Refine* SAM 6/3 0.85 1.95 1.28 3.17 9.04 5.28 7.49 14.4 4.14 7.43 5.51
3 3B Inter + Scribble Refine* SAM 6/3 0.8 4.11 2.95 2.19 3.13 3.36 7.14 9.85 4.02 4.67 4.22
4 3B Inter + Scribble Refine* SAM 6/3 0.71 2.45 2.01 2.21 2.28 1.89 6.08 9.09 3.22 3.57 3.35
5 3B Inter + Scribble Refine* SAM 6/3 0.76 1.99 2.46 2.51 2.16 1.42 6.15 9.34 3.19 3.47 3.34
0 3B Inter + Scribble Refine* SAM2 6/3 69.78 57.59 67.71 69.08 69.06 73.11 68.35 63.24 37.21 63.27 63.84
1 3B Inter + Scribble Refine* SAM2 6/3 10.03 20.17 25.37 19.13 23.65 18.12 25.37 27.66 11.34 15.93 19.68
2 3B Inter + Scribble Refine* SAM2 6/3 5.07 6.98 10.53 7.31 8.35 3.99 20.7 19.58 5.61 6.41 9.45
3 3B Inter + Scribble Refine* SAM2 6/3 4.23 4.43 7.8 6.93 5.64 3.72 14.43 16.35 4.89 5.34 7.38
4 3B Inter + Scribble Refine* SAM2 6/3 3.99 3.84 6.96 7.22 4.86 3.53 13.37 15.02 4.56 4.9 6.83
5 3B Inter + Scribble Refine* SAM2 6/3 3.41 3.7 6.75 5.7 3.99 2.79 12.49 13.59 4.3 4.47 6.12
0 3B Inter + Scribble Refine* SamMed 2D 6/3 27.83 42.22 52.7 60.66 44.76 38.63 55.31 55.81 33.46 38.66 45.0
1 3B Inter + Scribble Refine* SamMed 2D 6/3 28.4 44.72 57.53 63.33 45.28 41.42 61.61 59.73 40.44 38.8 48.13
2 3B Inter + Scribble Refine* SamMed 2D 6/3 24.29 42.41 53.02 59.24 40.27 35.69 62.86 58.85 41.23 34.93 45.28
3 3B Inter + Scribble Refine* SamMed 2D 6/3 20.91 39.69 47.48 52.73 35.55 28.2 62.08 54.89 40.02 31.59 41.31
4 3B Inter + Scribble Refine* SamMed 2D 6/3 18.76 37.29 41.46 46.5 31.73 23.98 58.46 51.86 38.07 28.74 37.69
5 3B Inter + Scribble Refine* SamMed 2D 6/3 17.07 35.06 37.39 41.62 29.1 20.8 53.9 49.27 36.27 26.66 34.71
@@
Table 5: Interactive refinement results for 2D models across 5 iterations. The initial prediction is made either using a single point per slice or one of our proposed prompting schemes. If the previous point prompt is not reused during refinement, indicated by a *, the performance drops. This also explains the drop in performance for the 3 Box interpolation after the initial prompt. The unrealistic slice-wise refinement (1 interaction per slice) is only slightly better than our proposed scribble refinement method (3 interactions).
@@
0 1 center PPV + 1 PPV Refine* SamMed 3D 1/1 2.03 12.15 24.11 26.9 15.09 19.64 72.66 53.03 50.4 26.51 30.25
1 1 center PPV + 1 PPV Refine* SamMed 3D 1/1 5.89 14.07 33.03 35.67 14.42 24.93 64.15 72.35 42.64 24.86 33.2
2 1 center PPV + 1 PPV Refine* SamMed 3D 1/1 5.67 13.49 29.52 28.92 14.34 21.4 64.29 62.44 42.03 25.16 30.73
3 1 center PPV + 1 PPV Refine* SamMed 3D 1/1 5.57 13.64 31.57 35.51 14.21 22.82 63.35 65.94 41.12 24.79 31.85
4 1 center PPV + 1 PPV Refine* SamMed 3D 1/1 5.81 13.51 31.44 31.08 14.11 21.38 65.28 63.38 40.68 24.82 31.15
5 1 center PPV + 1 PPV Refine* SamMed 3D 1/1 5.46 13.56 32.19 33.49 14.18 24.28 62.12 65.2 40.41 25.12 31.6
0 1 center PPV + 1 PPV Refine* SamMed 3D Turbo 1/1 5.12 27.34 46.07 34.46 15.9 46.38 82.98 59.36 63.83 26.98 40.84
1 1 center PPV + 1 PPV Refine* SamMed 3D Turbo 1/1 2.81 21.83 27.6 17.69 8.25 30.88 79.91 72.82 61.07 21.37 34.42
2 1 center PPV + 1 PPV Refine* SamMed 3D Turbo 1/1 3.57 22.72 35.57 22.05 8.89 36.31 77.77 72.2 62.7 23.37 36.52
3 1 center PPV + 1 PPV Refine* SamMed 3D Turbo 1/1 3.25 22.22 32.87 17.88 8.74 33.27 77.73 66.93 61.51 22.17 34.66
4 1 center PPV + 1 PPV Refine* SamMed 3D Turbo 1/1 3.6 22.93 34.4 16.26 8.78 34.84 80.01 72.09 62.35 23.35 35.86
5 1 center PPV + 1 PPV Refine* SamMed 3D Turbo 1/1 3.09 21.72 32.32 14.83 7.66 31.79 76.55 70.69 62.1 22.19 34.3
@@
Table 6: Interactive refinement results for 3D models over 5 iterations. The initial interaction always starts from a central point of the target object, and refinement is performed either by randomly sampling positive or negative points (1 interaction) or by selecting a point using the proposed scribble refinement method. Scribble drawing is counted as three interactions. In contrast to 2D models, including the previous point prompt did not improve the performance. It is indicated by a * if previous points are reused during refinement prompting.
Prompter Model Interactions D1 D2 D3 D4 D5 D6 D7 D10 D8a D8b Average Path. Average Org. Average
SAM NORM 3P Inter 3 0.84 3.6 3.49 2.22 1.67 1.03 8.28 4.56 18.98 6.37 5.1 3.41 9.05
SAM2 NORM 3P Inter 3 1.38 3.67 4.75 4.18 3.41 1.63 10.32 5.18 24.27 9.74 6.85 5.06 11.04
SamMed 2D NORM 3P Inter 3 11.61 29.08 28.0 34.33 26.0 10.8 36.29 23.21 40.98 30.62 27.09 25.38 31.09
SAM NORM 3B Inter 6 13.13 54.04 63.34 58.64 64.26 71.35 65.52 59.23 56.21 67.36 57.31 57.66 56.49
SAM2 NORM 3B Inter 6 69.76 57.59 67.71 69.08 69.06 73.11 68.35 63.26 57.43 69.05 66.44 69.45 59.43
SamMed 2D NORM 3B Inter 6 27.83 42.22 52.7 60.66 44.76 38.63 55.31 38.66 54.26 57.36 47.24 48.18 45.05
MedSam NORM 3B Inter 6 40.14 40.66 46.18 48.16 41.41 40.75 54.42 40.63 54.72 55.23 46.23 46.61 45.34
SamMed 3D NORM 1 center PPV 1 2.07 12.15 24.11 26.9 15.11 19.64 72.66 26.51 63.96 42.53 30.56 29.0 34.21
SamMed 3D Turbo NORM 1 center PPV 1 5.18 27.34 46.07 34.46 15.91 46.38 82.98 26.98 58.89 60.1 40.43 41.58 37.73
SegVol NORM 1 center PPV 1 9.96 24.91* 38.49 31.36 3.17 33.92 71.01 30.73 47.82 53.52 34.49 34.49 34.48
SegVol NORM 3D Box 3 0.55 37.17* 68.11 69.72 63.21 50.13 89.95 49.45 91.07 68.89 58.83 58.65 59.23
Table 7: Performance comparison for pathologies and organs. The table presents the average performance for all pathological datasets (D1, D3–D7, D8b) and all healthy organ datasets (D2, D8a, D19). Dataset D8 is split into two parts: D8a, which includes the liver, and D8b, which focuses on liver tumors as a pathological structure. Dataset D9, containing annotations of bone fractures, was excluded as it does not fit neatly into either category of healthy or pathological structures. * SegVol trained on CT images from the D2 HanSeg dataset, while we tested on MRI images from the same patient cohort.
Model Prompter Interactions Average MRI Average CT Average
MedSam 3B Inter 6 43.78 42.33 44.4
SAM 3B Inter 6 53.69 43.5 58.05
SAM2 3B Inter 6 63.84 65.02 63.33
SamMed 2D 3B Inter 6 45.0 40.91 46.76
SamMed 3D 3 center PPV 3 26.66 11.84 33.01
SamMed 3D Turbo 3 center PPV 3 41.94 25.15 49.13
SegVol 3 center PPV 3 43.32 30.9 48.64
SegVol 3D Box 3 58.04 35.28 67.8
Table 8: Averaged CT and MRI performance: The table shows the DSC for prompting schemes simulating a realistic Human Effort, without refinement, averaged over all MRI datasets (D1-3) and over all CT datasets (D4-10). The performance for models trained on medical data is highly dependent on the training data distribution: For example, SegVol was exclusively trained on CT data, leading to subpar performance on MRI datasets.